OpenStudy (anonymous):

does any body know 1/6-x/2=x-5/3??

OpenStudy (anonymous):

i am not sure how to get the answer for this alittle help?

OpenStudy (anonymous):

\[\frac{1}{6}=\frac{x}{2}=x-\frac{5}{3}\]?

OpenStudy (anonymous):

i meant \[\frac{1}{6}-\frac{x}{2}=x-\frac{5}{3}\]

OpenStudy (anonymous):

how did u get that answer??

OpenStudy (anonymous):

oops yes that how u write it

OpenStudy (anonymous):

if the fractions annoy you, multiply both sides (everything) by 6 to clear them. \[6\times(\frac{1}{6}-\frac{x}{2})=6\times (x-\frac{5}{3})\] \[1-3x=6x-10\] \[11=9x\] \[\frac{11}{9}=x\] check my algebra.

OpenStudy (anonymous):

yes the faction did annoy me... thanks for the answer {thumbs up}

OpenStudy (anonymous):

welcome